# Limit theorems for a strongly irreducible product of independent random matrices under optimal moment assumptions

### v2

@@ 4. Construction of a globally aligned extraction
Let $10<\alpha,\beta<$ 1, we have $(1+_{*}\mathcal{G}_{1-\alpha})^{*(1+_{*}\mathcal{G}_{1-\beta})}=(1+_{*}\mathcal{G}_{1-\alpha\beta})$.
@@ 4.1. Statement of the result and motivation
Let $(u_{n})_{n}$, $(s_{n})_{n}$ and $(w_{k})_{k}$ be as in Lemma 4.5. To all integer $n$, we associate a random number $i_{n}$, we write $i_{n}=1$ when there is an integer $k$ such that $\overline{w}_{2k}=n$ and 00 otherwise. We first show that $(i_{n})_{n}\sim\mathcal{B}_{\alpha}^{\otimes\mathbb{N}}$. For that, we first observe that by definition of $\mathcal{G}_{\alpha}$, we have:
@@
    The data of $(p_{k})_{k}$ is independent of the joint data of $(\tilde{s}_{k})_{k}$ and $(w_{k})_{k}$ and $(v_{k})_{k}$ and $(\widetilde{\gamma}^{\hat{v}}_{k})_{k\in\{0,1,2\}+4\mathbb{N}}$.
@@
[…]
@@
Let $E$ be a Euclidean, Hermitian or ultra-metric over a local field $\mathbb{K}$. Let $\Gamma=\mathrm{End}(E)$ or $\Gamma=\mathrm{GL}(E)$ and let $\nu$ be a probability measure on $\Gamma$. Let $\left(\gamma_{n}\right)\sim\nu^{\otimes\mathbb{N}}$. Let $\delta:\varepsilon\mapsto\varepsilon^{6}/48$. Let $m$, let $20<\alpha,\varepsilon\leq 1/$ 2, let $m\geq 1$ and let $\tilde{\nu}_{s}$ be as in Corollary 3.17.
@@
Point (5) states that the conditional distribution of $\gamma^{\check{p}}_{2k+1}$ with respect to $\left(\widetilde{\gamma}^{\check{p}}_{k^{\prime}}\right)_{k^{\prime}\neq 2k+1}$ is almost surely $\frac{1}{4}$-Schottky for $\mathbb{A}$. This comes from (11) again. By the Schottky property, we have $\tilde{\nu}_{s}(C_{k})\geq 4/6$. Therefore, the product of the normalized restriction of $\tilde{\nu}_{s}$ to $C_{k}$, is bounded above by $46\nu_{s}/$ 4, which is $\frac{1}{4}$-Schottky.
@@
Point (7) states that for all $A\subset\Gamma\setminus K$, we have for all $n$ and almost surely $\mathbb{P}(\gamma_{n}\in A\,|\,(p_{k})_{k\in\mathbb{N}})\leq 2\nu(A)$. First note that data of $(w_{k})_{k}$ determines whether $\gamma_{n}=u_{k}$ or $\gamma_{n}=s_{k}$ for all $n$. Hence, for all $n$, we have almost surely $\mathbb{P}(\gamma_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}})=\mathds{1}({s_{n}=\gamma_{n}})\mathbb{P}(s_{n}\in A)\leq\nu(A)/(1-\alpha)\leq 2\nu(A)$. On the set $(\gamma_{n}=u_{k})$, we have $\mathbb{P}(\gamma_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}})=0$ so $\mathbb{P}(\gamma_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}},(p_{k})_{k\in\mathbb{N}})=0$. On the set $(\gamma_{n}=s_{k})$, we have $\mathbb{P}(\gamma_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}},(p_{k})_{k\in\mathbb{N}})=\mathbb{P}(s_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}},(p_{k})_{k\in\mathbb{N}})=\mathbb{P}(s_{n}\in A)$ by the independence assumption (5).
Let un now prove point (8) using Lemmas (2.11) and 2.16. These points are purely geometric properties on the support of $\nu$. The proof relies on points (9) and (10) in Theorem 4.7, which are also almost sure properties (though not properties of the support because we do not assume $\mathbb{A}$ to be closed in general). Point (8) states that for all $i<j\leq k$, for all $f\mathbb{A}^{\varepsilon}\gamma^{\check{p}}_{i}$ and for all $\gamma^{\check{p}}_{k}\mathbb{A}^{\varepsilon}h$, we have $f\gamma^{\check{p}}_{i}\cdot\gamma^{\check{p}}_{j-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{j}\cdot\gamma^{\check{p}}_{k}h$. We will prove it by induction but first let us show how Lemma 2.16 applies.
@@
Let $k\geq 1$ and let $\gamma^{\check{p}}_{k}\mathbb{A}^{\frac{\varepsilon}{2}}h$. We show that then $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{k}h$. When $k$ is odd by point (9) in Theorem 4.7, we have $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\varepsilon}\gamma^{\check{p}}_{k}$ and $\sigma\left(\gamma^{\check{p}}_{k}\right)\leq\varepsilon^{6}/48\leq\varepsilon^{2}/8$. By Lemma 2.9, we have $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{k}h$. When $k$ is even, by Lemma 2.16, we have $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{k}h$. By induction, we have $\gamma^{\check{p}}_{j}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{j+1}\gamma^{\check{p}}_{k}h$ for all $j\leq k$.
@@ 4.3. Pivot algorithm
Let $\Gamma$ be a metric semi-group [^6] We call metric semi-group a metric space endowed with a continuous and associative product map. endowed with a measurable binary relation $\mathbb{A}$. Let $50<\rho<1/$ 5 and let $\nu_{s}$ be a probability distribution on $\Gamma$ that is $\rho$-Schottky for $\mathbb{A}$. Let $(\gamma_{n})\in\Gamma^{\mathbb{N}}$ and $(\tau_{n})\in[0,1]^{\mathbb{N}}$. To $(\gamma_{n})$ and $(\tau_{n})$ we associate a family $(p^{j}_{k})_{j,k\in\mathbb{N}}$ that we define by induction on $j$ by the following algorithm. For $j=0$, we set $p^{0}_{k}=1$ for all $k\in\mathbb{N}$. Let $j\geq 0$, we assume that we have constructed a sequence of odd integers $(p^{j}_{k})_{k\geq 0}$ such that that $p^{j}_{2k+1}=1$ for all $k$, such that there is an integer $m_{j}$ such that $\overline{p}^{j}_{2m_{j}+1}=2j+1$ and such that $p^{j}_{2k}=1$ for all $k>m_{j}$.
@@
We fix a metric semi-group $\Gamma$, endow it with its Borel $\sigma$-algebra $\mathcal{A}_{\Gamma}$ and fix a measurable binary relation $\mathbb{A}$. We fix a constant $50<\rho<1/$ 5 and a probability distribution $\nu_{s}$ on $\Gamma$ that is $\rho$-Schottky for $\mathbb{A}$. Let $\Omega:=\Gamma^{\mathbb{N}}\times[0,1]^{\mathbb{N}}$ and let $\mathcal{A}_{\Omega}$ be the infinite product of the Borel $\sigma$-algebras on $\Gamma$ and $[0,1]$. For all $n\in\mathbb{N}$, we write $\gamma_{n}:\Omega\to\Gamma$ for the map $((\gamma^{\prime}_{k})_{k},(\tau^{\prime}_{k})_{k})\mapsto\gamma^{\prime}_{n}$ and $\tau_{n}:\Omega\to\Gamma$ for the map $((\gamma^{\prime}_{k})_{k},(\tau^{\prime}_{k})_{k})\mapsto\tau^{\prime}_{n}$. Then for all $n\in\mathbb{N}$, the maps $\gamma_{n}$ and $\tau_{n}$ are continuous and therefore measurable. Note that the algorithm describes in Definition 4.9 associates a family $(p_{k}^{j})_{j,k\geq 0}$ to all pair of sequences $((\gamma^{\prime}_{k})_{k},(\tau^{\prime}_{k})_{k})\in\Omega$ so it describes a function $(p_{k}^{j})_{j,k\geq 0}:\Omega\to\mathbb{N}^{\mathbb{N}\times\mathbb{N}}$.
@@
[…]
@@ 5. Proof of the results
In this section, we use Theorem 1.6 to prove the other results of this paper. In this discussion, $(E,\|\cdot\|)$ is a Euclidean, Hermitian or ultra-metric vector space over a local field $(\mathbb{K},|\cdot|)$. Unless specified otherwise, $\nu$ is a strongly irreducible an proximal probability distribution over $\mathrm{End}(E)$. The constants $20<\varepsilon\leq 1/$ 2 and $m$ are as in Theorem 1.6. On a probability space $(\Omega,\mathbb{P})$, we define two random sequences $(\gamma_{n})_{n\geq 0}\sim\nu^{\otimes\mathbb{N}}$ and $\left(p_{n}\right)_{n\geq 0}$ that satisfy the conclusion of Theorem 1.6.
@@ 5.2. Law of large numbers and large deviations inequalities for the pivot extraction
Let $(\gamma_{n})_{n\geq 0}$ be a random sequence in $\mathrm{End}(E)$. Let $(p_{k})_{k\geq 0}$ be a random sequence of integers. Let $20<\varepsilon\leq 1/$ 2. Assume that points (2), (3) and (4) from Theorem 1.6 hold. Then there exists a constant $\lambda^{\prime}\in(0,+\infty]$ such that for all $i_{0}\geq 1$, we have almost surely:
@@
This is a consequence of Lemma 2.11 with $f=\gamma^{p}_{2i}\cdots\gamma^{p}_{2j}$, $g=\gamma^{p}_{2j+1}$ and $h=\gamma^{p}_{2j+2}\cdots\gamma^{p}_{2k}$. By (3), we have $\sigma(g)\leq\varepsilon^{6}/48$ and by (8), we have $f\mathbb{A}^{\frac{\varepsilon}{2}}g\mathbb{A}^{\frac{\varepsilon}{2}}h$ so $\sigma(fgh)\leq\sigma(f)\frac{\varepsilon^{6}}{48}\sigma(h)\frac{64}{\varepsilon^{4}}$ and $\varepsilon\leq\frac{1}{2}$ so $34\varepsilon^{2}/3\leq 1/$ 3, which proves (85).
By (2) and (4), the map $T:\Gamma^{\mathbb{N}}\to\Gamma^{\mathbb{N}}\,;\,(g_{k})_{k\geq 0}\mapsto(g_{k+2})_{k\geq 0}$ preserves the measure $\mu$ defined as $(\gamma^{p}_{k+2})_{k\geq 0}\sim\mu$. We claim that $\mu$ is in fact ergodic for $T$. Simply note that for all $i>1$, the data of $\left(\gamma^{p}_{k}\right)_{2\leq k\leq 2i}$ is independent of $(\gamma^{p}_{k})_{k\geq 2i+2}=T^{i}\left((\gamma^{p}_{k})_{k\geq 2}\right)$. Let $A^{\prime}\subset\Gamma^{\mathbb{N}}$ be $T$-invariant and let $A=\left((\gamma^{p}_{k})_{k\geq 2}\in A^{\prime}\right)$, that way $\nu(A^{\prime})=\mathbb{P}(A)$. We claim that $\mu(A^{\prime})\in\{0,1\}$, which is equivalent to saying that $A$ is independent of itself. Then for all $i>0$, we have $A=\left((\gamma^{p}_{k})_{k\geq 2}\in T^{-i}(A^{\prime})\right)=\left((\gamma^{p}_{k})_{k\geq 2i+2}\in A^{\prime}\right)$ Therefore for all $i>0$, $A$ is independent of the data of $\left(\gamma^{p}_{k}\right)_{2\leq k\leq 2i}$ is independent of $A$. Hence $A$ is independent of the joint data of $\left(\gamma^{p}_{k}\right)_{k\geq 2}$ and therefore of itself, which proves that $\mu$ is ergodic for $T$.
@@
(86) $-\log\sigma\left(\gamma^{p}_{2}\cdots\gamma^{p}_{2kn+r}\right)\leq\sum_{j=0}^{k-1}-\log\sigma(\gamma^{p}_{2jn+2}\cdots\gamma^{p}_{2(j+1)n}).$

### v3

@@ 4. Construction of a globally aligned extraction
Let $0<\alpha,\beta<1$, we have $(1+_{*}\mathcal{G}_{1-\alpha})^{*(1+_{*}\mathcal{G}_{1-\beta})}=(1+_{*}\mathcal{G}_{1-\alpha\beta})$.
@@ 4.1. Statement of the result and motivation
Let $(u_{n})_{n}$, $(s_{n})_{n}$ and $(w_{k})_{k}$ be as in Lemma 4.5. To all integer $n$, we associate a random number $i_{n}$, we write $i_{n}=1$ when there is an integer $k$ such that $\overline{w}_{2k}=n$ and 0 otherwise. We first show that $(i_{n})_{n}\sim\mathcal{B}_{\alpha}^{\otimes\mathbb{N}}$. For that, we first observe that by definition of $\mathcal{G}_{\alpha}$, we have:
@@
    The data of $(p_{k})_{k}$ is independent of the joint data of $(\tilde{u}_{k})_{k}$ and $(w_{k})_{k}$ and $(v_{k})_{k}$ and $(\widetilde{\gamma}^{\hat{v}}_{k})_{k\in\{0,1,2\}+4\mathbb{N}}$.
@@
On point that is really important to note is that the joint data of the random sequences $(p_{k})_{k}$, $(v_{k})_{k}$ and $(w_{k})_{k}$ is independent of the random sequence $(\tilde{u}_{k})_{k}$. If we moreover assume that $\tilde{\nu}_{s}$ has compact support, this fact allows us to bound the tail of the conditional distribution of $\gamma_{n}$ for a given $n\geq 0$ with respect to the joint data of $(p_{k})_{k}$, $(v_{k})_{k}$ and $(w_{k})_{k}$ by the tail distribution of the marginals of $\tilde{\kappa}$. For example, if we assume that $\mathbb{E}(N(\gamma_{0}))$ is finite for a given continuous and sub-additive map $N$ and write $B:=\max\{N(s_{k})\,|\,0\leq k<m,\;(s_{0},\dots,s_{m-1})\in\mathrm{supp}(\tilde{\nu}_{s})\}$, then we have a uniform bound over $\mathbb{E}(N(\gamma_{n})\,|\,(p_{k})_{k},(v_{k})_{k},(w_{k})_{k})\leq C$ with $C=\frac{1}{1-\alpha}\mathbb{E}(N(\gamma_{0}))+B$, which in turn gives a uniform bound over $\mathbb{E}\left(N(\gamma^{\check{p}}_{k})\right)\leq C\mathbb{E}(\check{p}_{k})$ for all $k\in\mathbb{N}$. This argument is the main ingredient in the proof of Theorem 1.13.
@@
Let $E$ be a Euclidean, Hermitian or ultra-metric over a local field $\mathbb{K}$. Let $\Gamma=\mathrm{End}(E)$ or $\Gamma=\mathrm{GL}(E)$ and let $\nu$ be a probability measure on $\Gamma$. Let $\left(\gamma_{n}\right)\sim\nu^{\otimes\mathbb{N}}$. Let $\delta:\varepsilon\mapsto\varepsilon^{6}/48$. Let $m$, let $0<\alpha,\varepsilon\leq 1/2$, let $m\geq 1$ and let $\tilde{\nu}_{s}$ be as in Corollary 3.17.
@@
Point (5) states that the conditional distribution of $\gamma^{\check{p}}_{2k+1}$ with respect to $\left(\widetilde{\gamma}^{\check{p}}_{k^{\prime}}\right)_{k^{\prime}\neq 2k+1}$ is almost surely $\frac{1}{4}$-Schottky for $\mathbb{A}$. This comes from (11) again. By the Schottky property, we have $\tilde{\nu}_{s}(C_{k})\geq 4/6$. Therefore, the product of the normalized restriction of $\tilde{\nu}_{s}$ to $C_{k}$, is bounded above by $6\nu_{s}/4$, which is $\frac{1}{4}$-Schottky.
@@
Point (7) states that for all $A\subset\Gamma\setminus K$, we have for all $n$ and almost surely $\mathbb{P}(\gamma_{n}\in A\,|\,(p_{k})_{k\in\mathbb{N}})\leq 2\nu(A)$. First note that data of $(w_{k})_{k}$ determines whether $\gamma_{n}=u_{k}$ or $\gamma_{n}=s_{k}$ for all $n$. Hence, for all $n$, we have almost surely $\mathbb{P}(\gamma_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}})=\mathds{1}({u_{n}=\gamma_{n}})\mathbb{P}(u_{n}\in A)\leq\nu(A)/(1-\alpha)\leq 2\nu(A)$. On the set $(\gamma_{n}=s_{k})$, we have $\gamma_{n}\in K$ almost surely so $\mathbb{P}(\gamma_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}},(p_{k})_{k\in\mathbb{N}})=0$. On the set $(\gamma_{n}=u_{k})$, we have $\mathbb{P}(\gamma_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}},(p_{k})_{k\in\mathbb{N}})=\mathbb{P}(u_{n}\in A\,|\,(w_{k})_{k\in\mathbb{N}},(p_{k})_{k\in\mathbb{N}})=\mathbb{P}(u_{n}\in A)$ by the independence assumption (5).
Let un now prove point (8) using Lemmas 2.11 and 2.16. These points are purely geometric properties on the support of $\nu$. The proof relies on points (9) and (10) in Theorem 4.7, which are also almost sure properties (though not properties of the support because we do not assume $\mathbb{A}$ to be closed in general). Point (8) states that for all $i<j\leq k$, for all $f\mathbb{A}^{\varepsilon}\gamma^{\check{p}}_{i}$ and for all $\gamma^{\check{p}}_{k}\mathbb{A}^{\varepsilon}h$, we have $f\gamma^{\check{p}}_{i}\cdots\gamma^{\check{p}}_{j-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{j}\cdots\gamma^{\check{p}}_{k}h$.We will prove it by induction but first let us show how Lemma 2.16 applies.
@@
Let $k\geq 1$ and let $\gamma^{\check{p}}_{k}\mathbb{A}^{\frac{\varepsilon}{2}}h$. We show that then $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{k}h$. When $k$ is odd by point (9) in Theorem 4.7, we have $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\varepsilon}\gamma^{\check{p}}_{k}$ and $\sigma\left(\gamma^{\check{p}}_{k}\right)\leq\varepsilon^{6}/48\leq\varepsilon^{2}/8$. By Lemma 2.9, we have $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{k}h$. When $k$ is even, by Lemma 2.16, we have $\gamma^{\check{p}}_{k-1}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{k}h$. By induction, we have $\gamma^{\check{p}}_{j}\mathbb{A}^{\frac{\varepsilon}{2}}\gamma^{\check{p}}_{j+1}\cdots\gamma^{\check{p}}_{k}h$ for all $j\leq k$.
@@ 4.3. Pivot algorithm
Let $\Gamma$ be a metric semi-group [^6] We call metric semi-group a metric space endowed with a continuous and associative product map. endowed with a measurable binary relation $\mathbb{A}$. Let $0<\rho<1/5$ and let $\nu_{s}$ be a probability distribution on $\Gamma$ that is $\rho$-Schottky for $\mathbb{A}$. Let $(\gamma_{n})\in\Gamma^{\mathbb{N}}$ and $(\tau_{n})\in[0,1]^{\mathbb{N}}$. To $(\gamma_{n})$ and $(\tau_{n})$ we associate a family $(p^{j}_{k})_{j,k\in\mathbb{N}}$ that we define by induction on $j$ by the following algorithm. For $j=0$, we set $p^{0}_{k}=1$ for all $k\in\mathbb{N}$. Let $j\geq 0$, we assume that we have constructed a sequence of odd integers $(p^{j}_{k})_{k\geq 0}$ such that that $p^{j}_{2k+1}=1$ for all $k$, such that there is an integer $m_{j}$ such that $\overline{p}^{j}_{2m_{j}+1}=2j+1$ and such that $p^{j}_{2k}=1$ for all $k>m_{j}$.
@@
We fix a metric semi-group $\Gamma$, endow it with its Borel $\sigma$-algebra $\mathcal{A}_{\Gamma}$ and fix a measurable binary relation $\mathbb{A}$. We fix a constant $0<\rho<1/5$ and a probability distribution $\nu_{s}$ on $\Gamma$ that is $\rho$-Schottky for $\mathbb{A}$. Let $\Omega:=\Gamma^{\mathbb{N}}\times[0,1]^{\mathbb{N}}$ and let $\mathcal{A}_{\Omega}$ be the infinite product of the Borel $\sigma$-algebras on $\Gamma$ and $[0,1]$. For all $n\in\mathbb{N}$, we write $\gamma_{n}:\Omega\to\Gamma$ for the map $((\gamma^{\prime}_{k})_{k},(\tau^{\prime}_{k})_{k})\mapsto\gamma^{\prime}_{n}$ and $\tau_{n}:\Omega\to\Gamma$ for the map $((\gamma^{\prime}_{k})_{k},(\tau^{\prime}_{k})_{k})\mapsto\tau^{\prime}_{n}$. Then for all $n\in\mathbb{N}$, the maps $\gamma_{n}$ and $\tau_{n}$ are continuous and therefore measurable. Note that the algorithm describes in Definition 4.9 associates a family $(p_{k}^{j})_{j,k\geq 0}$ to all pair of sequences $((\gamma^{\prime}_{k})_{k},(\tau^{\prime}_{k})_{k})\in\Omega$ so it describes a function $(p_{k}^{j})_{j,k\geq 0}:\Omega\to\mathbb{N}^{\mathbb{N}\times\mathbb{N}}$.
@@
$((\gamma_{2k})_{k\geq 0},(m_{j})_{j\geq 0})\sim\int_{x\in X}(\delta_{x}\otimes\kappa)d\mu_{X}(x)=\mu_{X}\otimes\kappa,$
@@ 5. Proof of the results
In this section, we use Theorem 1.6 to prove the other results of this paper. In this discussion, $(E,\|\cdot\|)$ is a Euclidean, Hermitian or ultra-metric vector space over a local field $(\mathbb{K},|\cdot|)$. Unless specified otherwise, $\nu$ is a strongly irreducible an proximal probability distribution over $\mathrm{End}(E)$. The constants $0<\varepsilon\leq 1/2$ and $m$ are as in Theorem 1.6. On a probability space $(\Omega,\mathbb{P})$, we define two random sequences $(\gamma_{n})_{n\geq 0}\sim\nu^{\otimes\mathbb{N}}$ and $\left(p_{n}\right)_{n\geq 0}$ that satisfy the conclusion of Theorem 1.6.
@@ 5.2. Law of large numbers and large deviations inequalities for the pivot extraction
Let $(\gamma_{n})_{n\geq 0}$ be a random sequence in $\mathrm{End}(E)$. Let $(p_{k})_{k\geq 0}$ be a random sequence of integers. Let $0<\varepsilon\leq 1/2$. Assume that points (2), (3) and (4) from Theorem 1.6 hold. Then there exists a constant $\lambda^{\prime}\in(0,+\infty]$ such that for all $i_{0}\geq 1$, we have almost surely:
@@
This is a consequence of Lemma 2.11 with $f=\gamma^{p}_{2i}\cdots\gamma^{p}_{2j}$, $g=\gamma^{p}_{2j+1}$ and $h=\gamma^{p}_{2j+2}\cdots\gamma^{p}_{2k}$. By (3), we have $\sigma(g)\leq\varepsilon^{6}/48$ and by (8), we have $f\mathbb{A}^{\frac{\varepsilon}{2}}g\mathbb{A}^{\frac{\varepsilon}{2}}h$ so $\sigma(fgh)\leq\sigma(f)\frac{\varepsilon^{6}}{48}\sigma(h)\frac{64}{\varepsilon^{4}}$ and $\varepsilon\leq\frac{1}{2}$ so $4\varepsilon^{2}/3\leq 1/3$, which proves (85).
By (2) and (4), the map $T:\Gamma^{\mathbb{N}}\to\Gamma^{\mathbb{N}}\,;\,(g_{k})_{k\geq 0}\mapsto(g_{k+2})_{k\geq 0}$ preserves the measure $\mu$ defined as $(\gamma^{p}_{k+2})_{k\geq 0}\sim\mu$. We claim that $\mu$ is in fact ergodic for $T$. Simply note that for all $i>1$, the data of $\left(\gamma^{p}_{k}\right)_{2\leq k\leq 2i}$ is independent of $(\gamma^{p}_{k})_{k\geq 2i+2}=T^{i}\left((\gamma^{p}_{k})_{k\geq 2}\right)$. Let $A^{\prime}\subset\Gamma^{\mathbb{N}}$ be $T$-invariant and let $A=\left((\gamma^{p}_{k})_{k\geq 2}\in A^{\prime}\right)$, that way $\nu(A^{\prime})=\mathbb{P}(A)$. We claim that $\mu(A^{\prime})\in\{0,1\}$, which is equivalent to saying that $A$ is independent of itself. Then for all $i>0$, we have $A=\left((\gamma^{p}_{k})_{k\geq 2}\in T^{-i}(A^{\prime})\right)=\left((\gamma^{p}_{k})_{k\geq 2i+2}\in A^{\prime}\right)$ Therefore for all $i>0$, AA is independent of the data of $\left(\gamma^{p}_{k}\right)_{2\leq k\leq 2i}$ is independent of $A$. Hence $A$ is independent of the joint data of $\left(\gamma^{p}_{k}\right)_{k\geq 2}$ and therefore of itself, which proves that $\mu$ is ergodic for $T$.
@@
(86) $\log\sigma\left(\gamma^{p}_{2}\cdots\gamma^{p}_{2kn+r}\right)\leq\sum_{j=0}^{k-1}-\log\sigma(\gamma^{p}_{2jn+2}\cdots\gamma^{p}_{2(j+1)n}).$